Suasan is three times as old as nancy, and Carrie is 5 years younger than Susan. Three years ago the sum of their ages was 56. What was nancy age four years ago?
Answer by stanbon(75887)

Susan is three times as old as nancy, and Carrie is 5 years younger than Susan. Three years ago the sum of their ages was 56. What was nancy age four years ago?
--------------
Equations:
s = 3n
c = s-5
s-3 + n-3 + c-3 = 56
-----------
s + n + c = 65
Substitute for n and for c; solve for "s"::
---------
s + (s/3) + s-5 = 65
-----------
s + (1/3)s + s = 65
-----
(7/3)s = 65
s = (3/7)65
s = 27.86 years (Susan's age now)
-----
n = (1/3)s = (1/3)27.86 = 9.29 years (Nancy's age now)
-----
c = s-5 = 22.86 years (Susan's age now)
==========================================
Nancy's age 4 years ago:: n-4 = 5.29 years
==========================================
Cheers,
Stan H.
-------------------